Required information Problem 10-2B Record equity transactions and indicate the effect on the balance sheet equation (LO10-2,View transaction ist Journal entry worksheet 2 3 4 5 6 Record the issuance of 150 additional shares of preferred stock for $2Journal entry worksheet Record the declaration of cash dividend on both common and preferred stock of $0.40 per share to allJournal entry worksheet < 1 2 3 5 6 Record the cash dividends declared on June 1. Note: Enter debits before credits Date GeneJournal entry worksheet < 1 2 3 4 5 Record the purchase of 250 shares of common treasury stock for $10 per share. Note: EnterJournal entry worksheet < 2 3 4 5 6 Record the resell 150 shares of treasury stock purchased on August 1 for $12 per share. NRequired information Problem 10-2B Record equity transactions and indicate the effect on the balance sheet equation (LO10-2,
Answer #1
Nautical
Journal entries
Date Account Debit Credit Calculation
1-Mar Cash $     16,900 =1,300*13
Common Stock $        1,300 =1,300*1
Additional Paid in Capital $     15,600
1-Apr Cash $        3,450 =150*23
Preferred Stock $        1,500 =150*10
Additional Paid in Capital $        1,950
1-Jun Common Dividends $        1,040 =(1,300+1,300)*0.40
Preferred Dividends $           120 =(150+150)*0.40
Dividends Payable $        1,160
30-Jun Dividends Payable $        1,160
Cash $        1,160
1-Aug Treasury Stock $        2,500 =250*10
Cash $        2,500
1-Oct Cash $        1,800 =150*12
Treasury Stock $        1,500 =150*10
Additional Paid in Capital $           300
Nautical
Transactions Total Assets Total Liabilties Total Stockholders Equity
Issue common stock + +
Issue preferred stock + +
Declare cash dividends + -
Pay cash dividends - -
Purchase treasury stock - -
Reissue treasury stock + +